Leo a working student from STEM, rides a bicycle from house to school to save money. If the mean life of a tire is 10,000 km and the standard deviation is 200 km. 68% of all tires will have a life between how many kilometers?


1
Expert's answer
2021-10-05T11:10:25-0400

Let X be a random variable representing the life of a tire. Then,

"X"~"N(\mu,\sigma^2)" where,

"\mu=10,000"

"\sigma^2=200^2=40,000"

Hence, "X"~"(10000, 40000)"

According to empirical rule, the area covered between mean minus 1 standard deviation and mean plus 1 standard deviation of a Normal distribution is 68%. This can be written as,

"p(\mu-\sigma\lt X \lt \mu+\sigma)=68\%"

"p(10000-200 \lt X \lt 10000+200)=68\%"

"p(9800 \lt X \lt 10200)=68\%"

Therefore, "68\%" of all tires will have a life between 9800km and 10200km.